OpenStudy (anonymous):

which expression is equivalent to (-2)(a+6) a. -2a+6 b. 2a+12 c.-2a-12 d.-2a+12 medal to who helps

OpenStudy (shamim):

Wt do u think abt ur option C

OpenStudy (shamim):

U hv to multiply a by -2

OpenStudy (anonymous):

frankly I don't understand it at all lol

OpenStudy (shamim):

N u hv to multiply 6 by -2

OpenStudy (shamim):

-2(a+6) =-2*a-2*6 =?

OpenStudy (anonymous):

thanks I got it

OpenStudy (shamim):

u r most welcome